OpenStudy (anonymous):

WHAT IS THE BEST ESTIMATE OF 22% OF 90 A.10 B.20 C.30 D.60

OpenStudy (anonymous):

b or c

OpenStudy (anonymous):

I WILL GIVE MEADEL AND FAN

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

B

OpenStudy (michele_laino):

Please you have to calculate: \[\frac{ 22 }{ 100 }\times 90=...\]

OpenStudy (anonymous):

yeah its B

OpenStudy (anonymous):

because 22% of 90 is 19.8 @oregon234

OpenStudy (goalieboy):

It is B because \[90\times .22=19.8\]

OpenStudy (anonymous):

edennoyes can u help me with this test answer all the questions and put the questions on here so i can see the here is the link https://www.connexus.com/content/chrome/online/lessonViewer.aspx?header=true&idCourse=33795&idUnit=24624&idLesson=288347&page=1&idWebuser=1229262&idSection=392775&self=True&menu=off&resizeWindow=true

OpenStudy (anonymous):

i need a log in

OpenStudy (anonymous):

did u look at it tell me answer number 17

OpenStudy (goalieboy):

@oregon234 did you get my answers

OpenStudy (goalieboy):

What?

OpenStudy (goalieboy):

@oregon234 what is going on here

OpenStudy (anonymous):

i clicked the link but i need a username and a password

OpenStudy (goalieboy):

here use mine